# A Review of Gravitational Memory and BMS Frame Fixing in Numerical Relativity

### v1

@@ VI.2 Detectability of Memory Effects
While Fig. 3 is useful in that it clearly illustrates the fact that the energy flux is the contribution responsible the memory exhibited by the gravitational wave strain, it does not provide the overall magnitude of this effect. This is because when the strain is evaluated at a point on the celestial sphere, e.g., if the strain were observed by a gravitational wave detector, one needs to consider the sum of the strain’s modes weighted by the spin weighted spherical harmonics. As a result, while the memory looks fairly prominent in the bottom left panel of Fig. 3, if one evaluates the strain at a point on the sky the memory can be noticeably suppressed. This is in part because the $(2,0)$ spin weight $-2$ spherical harmonic is
@@
so the memory is maximized for systems that, from the binary’s viewpoint, are viewed edge-on ($\theta=\pi/2$) and is minimized for systems that are viewed face-on ($\theta=0$). This fact is clearly highlighted in Fig. 2, which shows what the whole gravitational wave strain looks like for this binary black hole system when evaluated at the point on the sky with $(\theta,\phi)=(\pi/2,0)$. As can be seen, even for an ideal orientation (as well as an ideal binary, i.e., equal mass and large black hole spins in the direction of the orbital angular momentum), the net memory that is induced on the gravitational wave detector is $\lesssim 50\%$ of the magnitude of the full signal.
Furthermore, a unique challenge for observing memory in a real-world detector is the fact that detectors do not observe the GW strain in the time domain, but rather the frequency domain. This is an issue because the part of the strain that sources the memory, even though it can have a large amplitude, is a low-frequency effect since it resembles a step function and is not as oscillatory as the other modes of the strain. We highlight this in Fig. 4, which shows the Amplitude Spectral Density (ASD), i.e., the root of the Power Spectral Density (PSD), of the gravitational wave strain evaluated at $(\theta,\phi)=(\pi/2,0)$. We also show the contributions to the strain from the $(2,\pm 2)$ modes and the energy flux. The NR system is the same as that in Fig. 2. Before Fourier transforming the waveform we first window the waveform data using a Planck window with an $\epsilon$ value of $0.01$. As is shown, the ASD of the strain is primarily represented by the $(2,\pm 2)$ modes, while the energy flux, i.e., the memory, only contributes to the ASD at frequencies below 10Hz. [^20] Note that the reason why the ASD of the $(2,\pm 2)$ contribution falls off around 20Hz is because this is for a finite NR waveform whose initial $(2,2)$ frequency is 20Hz. If one examined the ASD for an infinitely long NR/PN hybrid, this falloff would not occur. This is what makes memory challenging to observe in current ground-based detectors. Because of seismic noise, LVK detectors are not sensitive to signals below $\sim$10Hz. In particular, the signal-to-noise ratio (SNR) $\rho$ of the strain and the memory contribution to the strain, i.e.,
@@
where $\tilde{h}(f)$ is the Fourier transform of the strain, $S_{n}(f)$ is the noise PSD, are $\approx$ 1,000 and $\approx$ 5. As a result, current efforts to observe memory rely on a procedure called “stacking” which combine the SNR estimates of the memory from various events to instead compute a type of population measurement Hübner et al. (2020, 2021); Grant and Nichols (2023). Put differently, these stacking efforts aim to show that memory has been detected in a population of events. To detect memory in a single event, we will most-certainly need to rely on future ground-based detectors like the Einstein Telescope or Cosmic Explorer (CE) or perhaps even one of the planned space-based detectors like LISA, which are more susceptible to the lower frequency regimes.
@@ VI.3 Memory Correction
Even so, Ref. Mitman et al. (2021b) showed that extrapolated waveforms can be corrected, through a sort of post-processing, to include the contribution from the memory and exhibit much better agreement with CCE waveforms. In Ref. Mitman et al. (2021b), it was found that for a wide range of binary simulations, the extrapolated strain waveforms simply seem to fail to capture the energy flux contribution in Eq. (21a). Thus, the authors argued that since the energy flux is only a function of the strain, the extrapolated strain waveforms can be self-consistently corrected to include the missing memory contribution. In doing so, they found that with such a correction the extrapolated waveforms then satisfy the supertranslation conservation law to a higher degree and better match the CCE waveforms. For completeness, we also show this phenomenon in Fig. 5. By eye, one can easily see that once the extrapolated strain waveform is corrected to include the missing energy flux contribution, then it agrees much better with the CCE waveform.
@@ VII BMS Frame Fixing
Up until this point, we have primarily been focused on utilizing the charge/flux perspective enlightened through the BMS group to understand the means by which the gravitational wave strain is sourced and memory effects. For the rest of this review, we will turn out attention to how the transformations of the BMS group are crucial for performing robust analyses with NR waveforms and building waveform models to test Einstein’s theory of GR with gravitational wave detector data.
@@ VII.1 Fixing the frame with BMS charges
As is the case in every field of physics, fixing the frame of the system that one is studying is vitally important to ensure that the observed phenomenon are physical and not just gauge artifacts. For binary black hole mergers, the situation is no different. In particular, whenever we study the gravitational radiation from a binary merger, we are often implicitly fixing part of the frame without knowing it. For example, by specifying a $\hat{z}$-axis, e.g., the direction of the binary’s orbital angular momentum or the direction of the remnant black hole’s spin axis, with respect to which we construct spherical harmonics to decompose the gravitational wave strain in, we are inherently fixing the rotation freedom by constructing a canonical direction. However, while this aspect of the frame fixing may seem trivial, there are other freedoms that need to be fixed that are more subtle.
@@
While at first this business of fixing the frame may seem trivial, we stress that performing this procedure is of the utmost importance to extract meaningful physics. To help highlight this, we present Fig. 6, which shows the consequence of not accounting for the difference in frames between an NR and a PN system. One important aspect of waveform modeling is constructing waveforms which span the entire frequency band of GW detectors. This means that one must build waveforms that contain thousands of orbits. NR simulations, however, typically only contain tens or perhaps hundreds of orbits, but not nearly as many as required to span the frequency bands of current and future observatories. Therefore, there is a constant need to “hybridize” finite NR waveforms with perturbative solutions that are accurate during the early inspiral regimes and can be more efficiently calculated. Often these hybrid waveforms are built by hybridizing finite NR waveforms with PN waveforms. This results in waveforms which are PN during the early inspiral phase, a smooth blend of PN and NR over a window called the hybridization window, and NR for the remainder of the binary coalescence. As is shown by Fig. 6, however, if one does not account for the freedoms resulting from the BMS group, i.e., if one does not map the NR system to be in the same BMS frame as the PN system, then this hybridization procedure will fail in the sense that it will introduce unphysical features in the hybrid waveform. Fortunately, this issue can be mitigated rather easily by simply mapping the NR system to be in the right frame. But, in order to do so, one needs to know how the various BMS symmetries transform the asymptotic coordinates as well as the asymptotic data of interest.
Fortunately, this was already presented in Sec. II.5. With both Eq. (7) and Eqs. (9), one has all of the information that is needed to transform asymptotic data. The only ingredient that remains for fixing the frame of the asymptotic data is which charges one should use to constrain the BMS freedoms. Normally, one could simply use the charges presented in Eqs. (23) and Eq. (25). However, because it is often the case that one wishes to compare NR systems to PN predictions for which there is only the strain and not the Weyl scalars, it turns out that there are more convenient charges that can be used. In particular, rather than working with the translation and boost charges, i.e., Eq. (23a) and Eq. (23c), it is instructive to instead construct the center-of-mass charge
@@
This charge is useful because we will almost always want to examine systems in their center-of-mass frame, which can be mapped to by finding the translation and boost which minimize Eq. (VII.1). As for the rotation charge, again because we typically do not have access to either of the $\Psi_{1}$ and $\Psi_{2}$ Weyl scalars in PN theory, one typically utilizes a rotation-like charge that can be computed from the strain. In particular, in Ref. Boyle (2013) such a rotation charge was built by finding the angular velocity which keeps the radiative fields, e.g., the strain, as constant as possible in the corotating frame of the binary system. This “charge” is
@@
When the system only consists of one black hole, as is the case, for example, when studying the ringdown phase, this prescription can break down. In this case, it is more useful to fix the rotation freedom using the intrinsic spin of the black hole, to ensure that one is in the frame of the individual black hole. To compute this, one can either map to the center-of-mass frame of the black hole and compute the angular momentum charge, or, for simplicity, they can instead compute
@@
is the velocity vector, and the vectors $\vec{J}$ and $\vec{K}$ are the angular momentum and boost charges found in Eq. (23b) and Eq. (23c). With this “charge”, one can then map to the frame of the black hole by, say, finding the rotation which aligns this charge with the positive $\hat{z}$-axis.
For fixing the supertranslation, since for fixing the Poincaré transformations we have been able to rely on the charges in Eqs. (23), one may naively think that we can simply use the supermomentum charge in Eq. (25). This charge, however, turns out to be supertranslation-invariant in nonradiative regimes of future null infinity, i.e., regimes where the news is zero. As a result, it cannot be used to fix the supertranslation freedom. [^21] This phenomenon should not necessarily come as a surprise, as this also happens for the translation-invariant momentum charge. Instead, one has to construct a different supertranslation “charge” which transforms in a meaningful way. In Ref. Moreschi (1986, 1988); Moreschi and Dain (1998); Dain and Moreschi (2000), such a charge was presented. It is simply Eq. (26) with $p=1$ and $q=0$ and is the Moreschi supermomentum:
@@
An important property of the Moreschi supermomentum is that, because it transforms as
@@
in Minkowski space where $m=0$, minimizing the Moreschi supermomentum corresponds to finding the supertranslation that ensures the spacetime is really Minkowski, not supertranslated Minkowski. Furthermore, this property is true so long as the $\ell\geq 2$ components of the mass aspect are zero, which is the same as there being no Geroch supermomentum. This is useful because isolated black holes cannot have Geroch supermomentum, so if one instead uses of the Moreschi supermomentum to fix the BMS freedom, they ensure that they are always working in the frame that is the most natural frame from the perspective of the individual black hole. We will call the BMS frame with no Moreschi supermomentum the superrest frame, as is it in some sense an extension of the notion of a rest frame. To map to such a frame, one can simply solve Eq. (VII.1) for $\mathcal{P}_{M}^{\prime}=-M_{B}$, which yields
@@
Furthermore, in Ref. Dain and Moreschi (2000), the authors proved that, for a condition on the energy flux, which is always obeyed in nonradiative regimes of future null infinity, Eq. (39) always has a regular solution. It was also shown that Eq. (39) can be solved iteratively. That is, if one wishes to find the supertranslation that maps a system to the superrest frame at some time $u_{0}$, they can evaluate the right-hand side of Eq. (39) at time $u=u_{0}$, solve for $\alpha$, evaluate the right-hand side at time $u=\alpha$, solve for a new $\alpha$, etc., until $\alpha$ converges to a solution. This is the underpinning of the frame fixing program. Specifically, by utilizing this fact, one can always solve for the infinite degree of freedom in the BMS transformation and map to a target BMS frame in an iterative fashion. For example, to map a NR system to the PN BMS frame, one can
@@
    Construct a window during the early inspiral phase over which to align an NR system to a PN system;
@@
For mapping to the frame of an individual black hole, like a remnant black hole, the process is identical, but in step (ii) one can minimize the Moreschi supermomentum, since the target Moreschi supermomentum is $-M_{B}$ and in step (iii) the rotation charge should be the spin of the black hole (see Eq. (33)). Furthermore, step (v) is no longer necessary since there is no preferred way to fix the time or phase freedom from the perspective of the remnant black hole. In the subsequent sections, we explain in more detail how to map a system to either the PN BMS frame or the superrest frame of an individual black hole and provide some numerical results showing why frame fixing is important and the success of this procedure at mapping NR systems to some reasonable BMS frame.

### v2

@@ VI.2 Detectability of Memory Effects
While Fig. 3 is useful in that it clearly illustrates the fact that the energy flux is the contribution responsible for the memory exhibited by the gravitational wave strain, it does not provide the overall magnitude of this effect. This is because when the strain is evaluated at a point on the celestial sphere, e.g., if the strain were observed by a gravitational wave detector, one needs to consider the sum of the strain’s modes weighted by the spin weighted spherical harmonics. As a result, while the memory looks fairly prominent in the bottom left panel of Fig. 3, if one evaluates the strain at a point on the sky the memory can be noticeably suppressed. This is in part because the $(2,0)$ spin weight $-2$ spherical harmonic is
@@
so the memory is maximized for systems that, from the binary’s viewpoint, are viewed edge-on ($\theta=\pi/2$) and is minimized for systems that are viewed face-on ($\theta=0$). This fact is clearly highlighted in Fig. 2, which shows what the whole gravitational wave strain looks like for a binary black hole system when evaluated at the point on the sky with $(\theta,\phi)=(\pi/2,0)$. As can be seen, even for an ideal orientation (as well as an ideal binary, i.e., equal mass and large black hole spins in the direction of the orbital angular momentum), the net memory that is induced on the gravitational wave detector is $\lesssim 50\%$ of the magnitude of the full signal.
Furthermore, a unique challenge for observing memory in a real-world detector is the fact that detectors do not analyze the GW strain in the time domain, but rather the frequency domain. This is an issue because the part of the strain that sources the memory, even though it can have a large amplitude, is a low-frequency effect since it resembles a step function and is not as oscillatory as the other modes of the strain. We highlight this in Fig. 4, which shows the Amplitude Spectral Density (ASD), i.e., the root of the Power Spectral Density (PSD), of the gravitational wave strain evaluated at $(\theta,\phi)=(\pi/2,0)$. We also show the contributions to the strain from the $(2,\pm 2)$ modes and the energy flux. The NR system is the same as that in Fig. 2. Before Fourier transforming the waveform we first pre-process it following Ref. Chen et al. (2024), i.e., we subtract a linear function from the waveform to remove the offset due to the memory effect and we apply a Planck window with an $\epsilon$ of $10^{-4}$ McKechan et al. (2010). As is shown, the ASD of the strain is primarily represented by the $(2,\pm 2)$ modes, while the energy flux, i.e., the memory, only contributes to the ASD at frequencies below 10Hz. This is what makes memory challenging to observe in current ground-based detectors. Because of seismic noise, LVK detectors are not sensitive to signals below $\sim$10Hz. In particular, the signal-to-noise ratio (SNR) $\rho$ of the strain and the memory contribution to the strain, i.e.,
@@
where $\tilde{h}(f)$ is the Fourier transform of the strain, $S_{n}(f)$ is the noise PSD, are $\approx$ 65 and $\approx$ 2. Because of this low SNR, current efforts to observe memory rely on a procedure called “stacking” which combine the SNR estimates of the memory from various events to instead compute a type of population measurement Hübner et al. (2020, 2021); Grant and Nichols (2023). Put differently, these stacking efforts aim to show that memory has been detected in a population of events. To detect memory in a single event, we will most-certainly need to rely on future ground-based detectors like the Einstein Telescope or Cosmic Explorer (CE) or perhaps even one of the planned space-based detectors like LISA, which are more susceptible to the lower frequency regimes.
@@ VI.3 Memory Correction
Even so, Ref. Mitman et al. (2021b) showed that extrapolated waveforms can be corrected, through a post-processing technique, to include the memory contribution and exhibit much better agreement with CCE waveforms. [^26] Ref. Talbot et al. (2018) also performed a correction to extrapolated waveforms to include memory, but did not compare the corrected waveforms to CCE waveforms, which naturally exhibit this effect. In Ref. Mitman et al. (2021b), it was found that for a wide range of binary simulations, the extrapolated strain waveforms simply seem to fail to capture the energy flux contribution in Eq. (22a). Thus, the authors argued that since the energy flux is only a function of the strain, the extrapolated strain waveforms can be self-consistently corrected to include the missing memory contribution. In doing so, they found that with such a correction the extrapolated waveforms then satisfy the supertranslation conservation law to a higher degree and better match the CCE waveforms. For completeness, we also show this phenomenon in Fig. 5. By eye, one can easily see that once the extrapolated strain waveform is corrected to include the missing energy flux contribution, it agrees much better with the CCE waveform.
Unfortunately, even with this useful memory correction, Ref. Mitman et al. (2021b) found that extrapolated waveforms still do not outperform CCE waveforms in terms of their accuracy and violation of the Bianchi identities. Consequently, future waveform models should use CCE waveforms.
@@ VII BMS Frame Fixing
Up until this point, we have primarily been focused on utilizing the charge/flux perspective enlightened through the BMS group to understand memory effects and the means by which the gravitational wave strain is sourced. For the rest of this review, we will turn our attention to how the transformations of the BMS group are crucial for performing robust analyses with NR waveforms and building waveform models to test Einstein’s theory of GR with gravitational wave detector data.
@@ VII.1 Fixing the frame with BMS charges
As is the case in every field of physics, fixing the frame of the system that one is studying is vitally important to ensure that the observed phenomena are physical and not just gauge artifacts. For binary black hole mergers, the situation is no different. In particular, whenever we study the gravitational radiation from a binary merger, we are often implicitly fixing part of the frame without knowing it. For example, by specifying a $\hat{z}$-axis, e.g., the direction of the binary’s orbital angular momentum or the direction of the remnant black hole’s spin axis, with respect to which we construct spherical harmonics to decompose the gravitational wave strain in, we are inherently fixing the rotation freedom by constructing a canonical direction. However, while this aspect of the frame fixing may seem trivial, there are other freedoms that need to be fixed that are more subtle.
@@
While at first this business of fixing the frame may seem trivial, we stress that performing this procedure is of the utmost importance to extract meaningful physics. To help highlight this, we present Fig. 6, which shows the consequence of not accounting for the difference in frames between a NR and a PN system. One important aspect of waveform modeling is constructing waveforms which span the entire frequency band of GW detectors. This means that one must build waveforms that contain thousands of orbits. NR simulations, however, typically only contain tens or perhaps hundreds of orbits, but not nearly as many as required to span the frequency bands of current and future observatories. Therefore, there is a constant need to “hybridize” finite NR waveforms with perturbative solutions that are accurate during the early inspiral regimes and can be more efficiently calculated. Often these hybrid waveforms are built by hybridizing finite NR waveforms with PN waveforms. This results in waveforms which are PN during the early inspiral phase, a smooth blend of PN and NR over a window called the hybridization window, and NR for the remainder of the binary coalescence. As is shown by Fig. 6, however, if one does not account for the freedoms resulting from the BMS group, i.e., if one does not map the NR system to be in the same BMS frame as the PN system, then this hybridization procedure will fail in the sense that it will introduce unphysical features in the hybrid waveform. Fortunately, this issue can be mitigated rather easily by simply mapping the NR system to be in the right frame. But, in order to do so, one needs to know how the various BMS symmetries transform the asymptotic coordinates as well as the asymptotic data of interest.
Fortunately, this was already presented in Sec. II.5. With both Eq. (7) and Eqs. (9), one has all of the information that is needed to transform asymptotic data. The only ingredient that remains for fixing the frame of the asymptotic data is which charges one should use to constrain the BMS freedoms. [^27] Note that throughout this review we ignore the subtle fact that the BMS charges presented in Eqs. (24) and in the remainder of this section are not supertranslation-invariant Christodoulou (1991); Chen et al. (2022); Javadinezhad et al. (2022); Javadinezhad and Porrati (2023, 2024). Normally, one could just use the charges presented in Eqs. (24) and Eq. (26). However, because it is often the case that one wishes to compare NR systems to PN predictions for which there is only the strain and not the Weyl scalars, it turns out that there are more convenient BMS charges that can be used. In particular, rather than working with the translation and boost charges, i.e., Eq. (24a) and Eq. (24c), it is useful to construct the center-of-mass charge Kozameh and Quiroga (2016); Compère et al. (2020)
@@
This measures the spacetime’s center-of-mass motion. [^28] One can think of the boost charge $K^{a}$ as measuring the translation away from the origin and the $uP^{a}$ term as measuring the change in center-of-mass due to a non-zero velocity. It is useful since we will always want to examine systems in their center-of-mass frame, which can be mapped to by finding the translation and boost which minimize Eq. (VII.1). As for the rotation charge, again because we typically do not have access to either of the $\Psi_{1}$ or $\Psi_{2}$ Weyl scalars in PN theory, one typically utilizes a rotation-like charge that can be computed from the strain. In particular, in Ref. Boyle (2013) such a rotation charge was built by finding the angular velocity which keeps the radiative fields, e.g., the strain, as constant as possible in the corotating frame of the binary system. This “charge” is
@@
When the system only consists of one black hole, as is the case, for example, when studying the ringdown phase, this prescription can break down. In this case, it is more useful to fix the rotation freedom using the intrinsic spin of the black hole, to ensure that one is in the frame of the individual black hole. To compute this, one can either map to the center-of-mass frame of the black hole and compute the angular momentum charge, or, for simplicity, they can instead compute Iozzo et al. (2021)
@@
is the velocity vector, and the vectors $\vec{J}$ and $\vec{K}$ are the angular momentum and boost charges found in Eq. (24b) and Eq. (24c). With this “charge”, one can then map to the frame of the black hole by, say, finding the rotation which aligns this charge with the positive $\hat{z}$-axis. Again, we stress that Eq. (33) and Eq. (35) have been introduced to help with fixing the rotation freedom by constructing a preferred axis, either through the angular velocity vector of a PN waveform or the spin of an isolated black hole.
For fixing the supertranslation, since for fixing the Poincaré transformations we have been able to rely on the charges in Eqs. (24), one may naively think that we can simply use the supermomentum charge in Eq. (26). This charge, however, turns out to be supertranslation-invariant in nonradiative regimes of future null infinity, i.e., regimes where the news is zero. As a result, it cannot be used to fix the supertranslation freedom.292929 This phenomenon should not necessarily come as a surprise, as this also happens for the translation-invariant momentum charge. Instead, one must construct a different supertranslation “charge” which transforms in a meaningful way. In Ref. Moreschi (1986, 1988); Moreschi and Dain (1998); Dain and Moreschi (2000), such a charge was presented. It is simply Eq. (27) with $p=1$ and $q=0$ and is the Moreschi supermomentum
@@
where $M_{\mathrm{ADM}}$ is the ADM mass Arnowitt et al. (1959). An important property of the Moreschi supermomentum is that, because it transforms as
@@
in Minkowski space where $m=0$, minimizing the Moreschi supermomentum corresponds to finding the supertranslation that ensures the spacetime is really Minkowski, not supertranslated Minkowski. Furthermore, this property is true so long as the $\ell\geq 2$ components of the mass aspect are zero, which is the same as there being no Geroch supermomentum. This is useful because isolated black holes cannot have Geroch supermomentum, so if one instead uses the Moreschi supermomentum to fix the BMS freedom, they ensure that they are always working in the frame that is the most natural frame from the perspective of the individual black hole. We will call the BMS frame in which the Moreschi supermomentum has zero-valued $\ell\geq 1$ modes the superrest frame, as it is in some sense an extension of the notion of a rest frame. To map to such a frame, one can simply solve Eq. (VII.1) for $\mathcal{P}_{M}^{\prime}=-M_{B}$, which yields
@@
Furthermore, in Ref. Dain and Moreschi (2000), the authors proved that, for a condition on the energy flux, which is always obeyed in nonradiative regimes of future null infinity, Eq. (41) always has a regular solution. It was also shown that Eq. (41) can be solved iteratively. That is, if one wishes to find the supertranslation that maps a system to the superrest frame at some time $u_{0}$, they can evaluate the right-hand side of Eq. (41) at time $u=u_{0}$, solve for $\alpha$, evaluate the right-hand side at time $u=\alpha$, solve for a new $\alpha$, etc., until $\alpha$ converges to a solution. This is the underpinning of the frame fixing program. Specifically, by utilizing this fact, one can always solve for the infinite degrees of freedom in the BMS transformation and map to a target BMS frame in an iterative fashion. For example, to map a NR system to the PN BMS frame, one can
@@
    Construct a window during the early inspiral phase over which to align a NR system to a PN system;
@@
For mapping to the frame of an individual black hole, like a remnant black hole, the process is identical, but in step (ii) one can minimize the Moreschi supermomentum, since the target Moreschi supermomentum is $-M_{B}$ and in step (iii) the rotation charge should instead be the spin of the black hole (see Eq. (35)). Also, step (v) is no longer necessary since there is no preferred way to fix the time or phase freedom from the perspective of the remnant black hole. In the subsequent sections, we explain in more detail how to map a system to either the PN BMS frame or the superrest frame of an individual black hole and provide some numerical results showing why frame fixing is important and the success of this procedure at mapping NR systems to some reasonable BMS frame.